QUESTION 29 YearUnits of Outpu Price Per Unit G4 20 25 30 3 me an economy that is producing only one product. Output and price data for a three-year period are shown in the table. If year 2 is chosen for the base year, in year 3 nominal GDP and real GDP, respectively, are $120 and $100 $30 and $5. 。$180 and $120 O $180 and $30.
Answer #1

Answer : 29) The answer is option C.

Nominal GDP for year 3 = Price of current year × Output of current year = 6 × 30 = $180

Real GDP for year 3 = Price of base year × Output of current year = 4 × 30 = $120

Therefore, option C is correct.